The lowest pressure attainable using the best available vacuum techniques is about 10−12N/m2.

1. At such a pressure, how many molecules are there per cm3 at 11 ∘C?

Express your answer using two significant figures.
Answer #1

let the numner of moles per cm^3 is n

P = 1 *10^-12 Pa

Using ideal gas equation

P * V = n * R * T

10^-12 * 10^-6 = n * 8.314 * (11 + 273)

solving for n

n = 4.235 *10^-22

number of molecules per cm^3 = NA * n

number of molecules per cm^3 = 4.235 *10^-22 * 6.022 *10^23

number of molecules per cm^3 = 255

the number of molecules per cm^3 is 255